OpenStudy (raheelafzaal):

plz help me d/dx(cot^-1)=?

OpenStudy (anonymous):

y = 1 / ( cscx + cot x ) = csc x - cot x ....... (1) dy/dx = d/dx( csc x ) - d/dx( cot x ) . . . . .= -csc x cot x - ( - csc² x ) . . . . .= csc x ( csc x - cot x ) ................... Ans. OR . .= y· csc x .......... from (1) .............. Ans. ___________________________________ Happy To Help ! ___________________________________

OpenStudy (experimentx):

the derivative of constant is zero.

OpenStudy (anonymous):

so my ans.. is correct

OpenStudy (experimentx):

not really, http://www.math.com/tables/derivatives/tableof.htm if it was cot^-1x|x=1, then answer would be -1/2